OpenStudy (anonymous):

how to write 33 1/3 percent in simplest form?

OpenStudy (tkhunny):

Eyeball problem. You should just recognize this one. 33 1/3 percent =1/3 -- Done!

OpenStudy (anonymous):

Why/how?

OpenStudy (anonymous):

100/3

OpenStudy (tkhunny):

2's 1/2 = 50% 2/2 = 100% 3's 1/3 = 33 1/3% 2/3 = 66 2/3% 4's 1/4 = 25% 2/4 = 50% 3/4 = 75% 5's 1/5 = 20% 2/5 = 40% 3/5 = 60% 4/5 = 80% Seriously, some you should just KNOW. No need to fiddle with them.

OpenStudy (anonymous):

(33*3+1)/3=100/3

OpenStudy (tkhunny):

100/3 is no good. It should be 100/3 %, or just 1/3.

OpenStudy (tkhunny):

6's 1/6 = 16 2/3% 2/6 = 33 1/3% 3/6 = 50% 4/6 = 66 2/3% 5/6 = 83 1/3% 7's are fun. I'll leave them for your private exploration.